# Efficient Inference in First Passage Time Models

### v1

@@ Abstract
First passage time models describe the time it takes for a random process to exit a region of interest and are widely used across various scientific fields. Fast and accurate numerical methods for computing the likelihood function in these models are essential for efficient statistical inference. Specifically, in mathematical psychology, generalized drift diffusion models (GDDMs) are an important class of first passage time models that describe the latent psychological processes underlying simple decision-making scenarios. GDDMs model the joint distribution over choices and response times as the first hitting time of a one-dimensional stochastic differential equation (SDE) to possibly time-varying upper and lower boundaries. They are widely applied to extract parameters associated with distinct cognitive and neural mechanisms. However, current likelihood computation methods struggle with common scenarios where drift rates covary dynamically with exogenous covariates in each trial, such as in the attentional drift diffusion model (aDDM). In this work, we propose a fast and flexible algorithm for computing the likelihood function of GDDMs based on a large class of SDEs satisfying the Cherkasov condition. Our method divides each trial into discrete stages, employs fast analytical results to compute stage-wise densities, and integrates these to compute the overall trial-wise likelihood. Numerical examples demonstrate that our method not only yields accurate likelihood evaluations for efficient statistical inference, but also significantly outperforms existing approaches in terms of speed.
Keywords— first passage time, first hitting time, drift diffusion model, attention, likelihood-based inference, numerical methods, Cherkasov condition
@@ 1 Introduction
First passage time models (also known as first hitting time models) are a fundamental class of stochastic process models that describe the time it takes for a random process to exit a region of interest. These models are widely used in various fields, including physics, finance, biology, and cognitive science, to analyze phenomena such as the time until a stock hits a particular price level, the time it takes for a neuron to fire or the time to reach a decision barrier that triggers a choice in psychological experiments [5, 39, 22]. Accurately computing the density of the first passage time is crucial for the practical application of these models, such as in scientific hypothesis testing and parameter estimation. Since analytical solutions for many complex real-world first passage time models are unavailable, effective numerical methods are an indispensable part of the computational toolkit for applying these models. In this paper, we introduce a fast and flexible method for computing the first passage time density in a large class of models arising in cognitive neuroscience, mathematical psychology, and economics.
Consider a psychological experiment in which a subject must choose between two alternatives, A and B. The experimenter records both the choice and the time taken to make it. A common approach to modeling this behavior is through a stochastic process $X(t)$ that starts at zero and evolves randomly over time until it first reaches ${+}a$ or ${-}a$, where $a$ is the evidence threshold. The choice (A vs. B) is modeled by which boundary is hit (${+}a$ vs. ${-}a$) and the response time is modeled by the first hitting time.
@@
This basic DDM can be modified in a variety of ways to better model experimental data and test theories about the biological processes underlying decision making. For instance, the boundaries (±$\pm a$) can be asymmetric and vary over time to reflect changing pressures to make a decision; the starting position $X(0)$ can be nonzero or random, representing an initial decision bias; the rates of drifting towards the boundaries can depend on the choices and vary over time; and the random process $X(t)$ can have more complicated stochastic dynamics than Brownian motion [17]. We adopt the terminology from [57] and call these models generalized drift diffusion models (GDDMs). Simple, computationally tractable versions of GDDMs have found widespread use in cognitive neuroscience and mathematical psychology [49, 50, 21, 15, 72]. While more complex versions of GDDMs have been proposed to encode a broader range of scientific hypotheses, their quantitative assessment and applications are currently limited due to computational challenges in parameter estimation [17]. In particular, there do not exist fast algorithms for calculating the likelihood function of the parameters (i.e., the probability density of the hitting time on the boundary reached) in some of the more complex GDDMs that are currently of scientific interest.
@@
Computational speed is crucial in many cases for the following reason: The precise specification of the GDDM might change every time a subject makes a choice. An experiment with, say, $50$ subjects, each completing $200$ trials, might require evaluating the likelihoods of $10^{4}$ different versions of the GDDM. For instance, in equation (1) the drift rate $\mu$ might be different on every trial if the choices presented to subjects are different on every trial. The problem becomes much more difficult when the trial-specific drift rate can vary dynamically within a trial. Our original motivation to study this problem came from attentional drift diffusion models (aDDMs) [34], which have been proposed as a model for investigating how visual attention influences decision making. In aDDMs, the drift rate $\mu(t)$ in each trial is allowed to covary on a moment-by-moment basis, depending on the subject’s eye movements toward one choice option or another, which are also recorded in the experimental setup. The entire shape of $\mu(t)$ changes dramatically on every trial. This model is a generic example of the general class of GDDMs where instantaneous within-trial dynamics can be based on exogenous covariates such as brain activity or eye gaze position, hence complicating the numerical procedures for computing the first passage time density and further applications. Furthermore, hierarchical Bayesian inference is commonly used in cognitive science to estimate both group- and subject-level parameters simultaneously [70, 17]. The reliance on Markov chain Monte Carlo (MCMC) methods in this context further amplifies the demand for fast likelihood evaluations.
@@
for possibly random $X(0)$ and for functions $\mu$ and $\sigma$ satisfying the so-called Cherkasov condition [10, 53]. Loosely speaking, the Cherkasov condition means that the complicated SDE in equation (2) can be transformed into a simpler SDE more akin to equation (1). A special case that includes many cases of practical interest, such as the aDDM, is when $\mu$ and $\sigma$ depend only on $t$ and not on $X(t)$ [^1] In the cognitive science context, the terminology “DDM” is typically adopted to denote these models that directly integrate noisy accumulating evidence. The more general model in equation (2), including some examples shown later in this paper (e.g., the Ornstein-Uhlenbeck model), fall within the GDDM framework.. There exist numerical approaches and software for approximating hitting time densities in models even more general than equation (2), but in many cases they are currently too slow for practical use in real-world psychological experiments.
Our approach is closely related to the methods in [67, 42, 47, 68, 60] which approximate the cumulative distribution function of $X(T)$ at a fixed time $T$ given that the process has not reached the boundary before $T$. Both their method and ours transform the problem into a canonical representation, subdivide time into discrete stages that allow for analytic approximations, compute the final approximation through integration over these stages, and then map the result back to the original problem. However, our work is distinct in that we focus on computing the density of the hitting time rather than the distribution of the process at a fixed time. Additionally, our approach is significantly more computationally efficient, as we pay careful attention to computational details, such as the method of numerical integration. Each of these innovations is crucial for practical likelihood-based inference in modern experimental settings.
@@
The remainder of this article is structured in the following way. Section 2 begins with a mathematical description of the GDDM. Section 3 summarizes previous work related to this paper, including other efforts to address the computational challenges in first passage time models. In Section 4, we introduce our proposed method for computing the FPTDs of certain GDDMs, starting with analytical results for simple DDMs, then combining them sequentially based on Markovian properties. We also apply variable transformations and piecewise linear approximations to extend the method to more complex GDDMs, including those that allow for position-dependent leak and dynamic decision boundaries [64, 17]. Section 5 shows several numerical examples that demonstrate the speed and accuracy of our method for both likelihood computation and statistical inference tasks. Section 6 summarizes the conclusion of the present work, including its applicability and suggestions for future work. It is worth mentioning that while our examples are drawn from cognitive science and mathematical psychology, the method we propose is equally applicable to first passage time models in other fields. For instance, it could be extended to financial applications such as pricing double barrier options [36], as well as other areas where first passage times play a critical role.
@@ 2 Drift Diffusion Model
where $\mathbbm{1}_{E}(\omega)$ is the indicator function of $E$ that is $1$ if $\omega\in E$ and 00 otherwise.
If we consider consider the possible truncation of the process at $t=T<\infty$, the process can either exit through the upper boundary $u(t)$ before $T$, exit through the lower boundary $\ell(t)$ before $T$, or fail to hit either boundary by $T$ and instead reach the truncation boundary $t=T$, resulting in three sub-probability densities:
@@
As mentioned in the Introduction, we are motivated by practical problems for which we need to evaluate many different joint densities, each at perhaps a single point. Define $\boldsymbol{\Xi}=\{\mu(\cdot),\sigma(\cdot),u(\cdot),\ell(\cdot),q_{0}(\cdot)\}$ to be the configuration of a GDDM, where $q_{0}$ is a representation for the distribution of $X(0)$, usually a probability density function or a point mass at a deterministic starting point $x_{0}$. In statistical applications, $\boldsymbol{\Xi}$ will depend on unknown parameters and known covariates, some of which might vary with each trial. If the $i$-th trial results in the observation $(\tau_{i},C_{i})$, then the contribution to the likelihood for that trial is $f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i})$. This is simply the FPTD on the boundary $C_{i}$ evaluated at time $\tau_{i}$ using the GDDM configuration $\boldsymbol{\Xi}_{i}$ for the $i$-th trial. In experiments with a time limit, some trials might result in non-response, in which case the contribution to the likelihood is the NPP for that GDDM configuration, say $Q_{\boldsymbol{\Xi}_{i}}$. A usual experiment in the psychological and cognitive sciences, where GDDMs are ubiquitously applied, produces a dataset with thousands or tens of thousands of trials (across multiple subjects). Consequently, fast and accurate methods for evaluating FPTDs and NPPs are crucial for the practical use of likelihood-based inference in these models.
@@ 3 Related Work
Analytical formulas. For certain simple DDMs, analytical formulas for the first passage time densities or distributions can be derived, often in the form of infinite series. These methods are fast and accurate, but are typically applicable only for very specific cases. For example, they are available for the standard DDM where all parameters $\mu,\sigma,x_{0},u,\ell$ are constant [41, 6, 26]. The approach in this paper is based on numerically extending these fast special cases to much broader classes of GDDMs. We note that the methods mentioned in the Section 1 for computing distributions [67, 42, 47, 68] can also be used to derive densities by numerically differentiating the distribution. However, the use of finite difference will introduce additional numerical error.
@@
Numerical solutions of governing equations. These methods exploit the known connections between diffusion processes and partial differential equations (PDEs), called Kolmogorov equations (see the Chapter 8 of [69] for an introduction, and Chapter 5.7 of [32] for a mathematically rigorous treatment). Both the Kolmogorov forward equation (KFE, also known as Fokker-Planck equation in physics) [57, 13] and the Kolmogorov backward equation (KBE) [66, 65] have been exploited to derive the FPTDs. Moreover, the first passage time problem can be formulated into Volterra integral equations of renewal type [58, 46], where applications to DDMs [59, 43] have also been explored. In these works, numerical methods, particularly finite difference methods [62] due to their simplicity, have been developed to solve these equations, enabling the efficient computation of FPTDs in certain cases. Other numerical methods like discontinuous Galerkin methods have also been considered, see [45] and references therein. However, these approaches remain computationally expensive, especially in settings where the GDDM configuration differs across trials, including but not limited to cases with trial-specific covariates that vary dynamically over time. For a comprehensive review of these methods, we refer the readers to [54].
Monte Carlo methods. Monte Carlo methods [24, 31, 25], including exact simulation methods [29, 30], generally require large sample sizes to ensure accurate likelihood approximations, making them slow in practice. However, we apply Monte Carlo methods in Section 5 to verify the results of our approach. The simulation results can support statistical inference by enabling the comparison of summary statistics between observed and simulated data [51]. Methods such as density estimation [14] and approximate Bayesian computation (ABC) [55, 3] are commonly employed for this purpose.
Simulation Based Inference (SBI). Another alternative approach, which is gaining traction in recent years, is to train machine learning models on simulated data as a surrogate for likelihoods. The surrogate likelihood function is commonly represented by a neural network [44, 18, 48, 7, 27] or a Gaussian process [1, 71]. This allows the expensive Monte Carlo simulations to be carried out prior to data collection. The inexpensive amortized likelihoods can be deployed at inference time to speed up parameter estimation [11, 18]. However, for cases where $\boldsymbol{\Xi}$ has high-dimensional parameters or covariates that dynamically vary in time, the amortization becomes cumbersome or intractable. Another current drawback of SBI is that the user has less control over the quality of the resulting approximations. We are not aware of any SBI work specifically applied to the specific models highlighted in this paper. However, we mention SBI as a promising general purpose alternative, which in principle can have bearing on the model domain under discussion in this work.
@@ 4.1 Single-stage Model
[28] derives the following analytical formulas [^3] Another work along this direction is [2]. We adopt the results in [28] since they are simpler and were shown to agree with those of [2]. Equation (9), (10), and (11) have been slightly rewritten for use in this paper, compared to their original form in [28]. for FPTDs and NPDs in the single-stage model for the special case when $x_{0}=0$ and $\sigma=1$: We define $\overline{a}=\frac{a_{1}+a_{2}}{2},\ \overline{b}=\frac{b_{1}+b_{2}}{2},\ c=a_{1}-a_{2},\ b=\frac{b_{2}-b_{1}}{2}$, then
@@
    $\quad\frac{1}{\sqrt{2\pi t^{3}}}e^{-\frac{b}{c}a_{1}^{2}+a_{1}\delta_{u}-\frac{1}{2}\delta_{u}^{2}t}\sum_{j=0}^{\infty}(-1)^{j}\alpha_{j}e^{\left(\frac{b}{c}-\frac{1}{2t}\right)\alpha_{j}^{2}}\mathbbm{1}_{(0,T]}(t)$
@@
    $\quad\frac{1}{\sqrt{2\pi t^{3}}}e^{-\frac{b}{c}a_{2}^{2}-a_{2}\delta_{\ell}-\frac{1}{2}\delta_{\ell}^{2}t}\sum_{j=0}^{\infty}(-1)^{j}\beta_{j}e^{\left(\frac{b}{c}-\frac{1}{2t}\right)\beta_{j}^{2}}\mathbbm{1}_{(0,T]}(t)$
@@
    $\quad\frac{1}{\sqrt{2\pi T}}e^{(\mu-\overline{b})x-\frac{1}{2}(\mu^{2}-\overline{b}^{2})T}\bigg{\{}e^{-\frac{y^{2}}{2T}}+\sum_{j=1}^{\infty}\bigg{[}e^{4bj(jc-\overline{a})-\frac{(y-2jc)^{2}}{2T}}$
    $\quad-e^{2b(2j-1)\left(jc-a_{1}\right)-\frac{(y+2jc-2a_{1})^{2}}{2T}}+e^{4bj(jc+\overline{a})-\frac{(y+2jc)^{2}}{2T}}$
    $\quad-e^{2b(2j-1)\left(jc+a_{2}\right)-\frac{(y-2jc-2a_{2})^{2}}{2T}}\bigg{]}\bigg{\}}\mathbbm{1}_{(\ell(T),u(T))}(x)$
@@
$f_{u}^{\text{single}}(t;\mu,\sigma,\mathcal{B},x_{0})=f_{u}^{\text{basic}}\Big{(}t;\frac{\mu}{\sigma},\frac{a_{1}-x_{0}}{\sigma},\frac{b_{1}}{\sigma},\frac{a_{2}-x_{0}}{\sigma},\frac{b_{2}}{\sigma},T\Big{)}$ (9)
$f_{\ell}^{\text{single}}(t;\mu,\sigma,\mathcal{B},x_{0})=f_{\ell}^{\text{basic}}\Big{(}t;\frac{\mu}{\sigma},\frac{a_{1}-x_{0}}{\sigma},\frac{b_{1}}{\sigma},\frac{a_{2}-x_{0}}{\sigma},\frac{b_{2}}{\sigma},T\Big{)}$ (10)
$q^{\text{single}}(x;\mu,\sigma,\mathcal{B},x_{0})=\frac{1}{\sigma}q^{\text{basic}}\Big{(}\frac{x-x_{0}}{\sigma};\frac{\mu}{\sigma},\frac{a_{1}-x_{0}}{\sigma},\frac{b_{1}}{\sigma},\frac{a_{2}-x_{0}}{\sigma},\frac{b_{2}}{\sigma},T\Big{)}$ (11)
@@
Figure 2: (a) A single-stage model is a GDDM with constant drift $\mu$ and two linear boundaries $u(t),\ell(t)$ starting at $x=x_{0}$ and truncated at $t=T$. The process can hit one of $u(t)$ and $\ell(t)$ before $T$, or hit neither. This results in three distinct subprobability distributions, corresponding to these three outcomes, with their subprobability densities plotted by red, purple, and green curves along the boundaries. Sample trajectories, colored to match their respective outcomes, are also displayed. The blue line represents the expected value of the process, given by $\mathbb{E}[X(t)]=x_{0}+\mu t$. In a single-stage model the total area under the three subprobability density curves sums to 1. (b) The integrated model extends the single-stage model by introducing random variability in the initial position $X(0)$. The expected value of the process in this model is given by $\mathbb{E}[X(t)]=\mathbb{E}[X(0)]+\mu t$, represented by the blue line. The density of $X(0)$, denoted by $q_{0}$, is illustrated by the grey curve. We allow $q_{0}$ to be a subprobability density function. In an integrated single-stage model, the total area under the three outcome subprobability density curves sums to the area under the initial grey curve.

### v2

@@ Abstract
First passage time models describe the time it takes for a random process to exit a region of interest and are widely used across various scientific fields. Fast and accurate numerical methods for computing the likelihood function in these models are essential for efficient statistical inference of model parameters. Specifically, in computational cognitive neuroscience, generalized drift diffusion models (GDDMs) are an important class of first passage time models that describe the latent psychological processes underlying simple decision-making scenarios. GDDMs model the joint distribution over choices and response times as the first hitting time of a one-dimensional stochastic differential equation (SDE) to possibly time-varying upper and lower boundaries. They are widely applied to extract parameters associated with distinct cognitive and neural mechanisms. However, current likelihood computation methods struggle in common application scenarios in which drift rates dynamically vary within trials as a function of exogenous covariates (e.g., brain activity in specific regions or visual fixations). In this work, we propose a fast and flexible algorithm for computing the likelihood function of GDDMs based on a large class of SDEs satisfying the Cherkasov condition. Our method divides each trial into discrete stages, employs fast analytical results to compute stage-wise densities, and integrates these to compute the overall trial-wise likelihood. Numerical examples demonstrate that our method not only yields accurate likelihood evaluations for efficient statistical inference, but also considerably outperforms existing approaches in terms of speed.
keywords:
first passage time, drift diffusion model, attention, likelihood-based inference, numerical methods, Cherkasov condition
@@ 1 Introduction
First passage time models (also known as first hitting time models) are a fundamental class of stochastic process models that describe the time it takes for a random process to exit a region of interest. These models are widely used in various fields, including physics, finance, biology, and cognitive science, to analyze phenomena such as the time until a stock hits a particular price level, the time it takes for a neuron to fire, a survival time of a transplant patient, or the time to reach a decision barrier that triggers a choice in psychological experiments [3, 36, 18, 32]. Accurately computing the density of the first passage time is crucial for the practical application of these models, such as in scientific hypothesis testing and parameter estimation. Since analytical solutions for many complex real-world first passage time models are unavailable, effective numerical methods are an indispensable part of the computational toolkit for applying these models. In this paper, we introduce a fast and flexible method for computing the first passage time density in a large class of models arising in cognitive neuroscience, mathematical psychology, and economics.
Consider a psychological experiment in which a subject must choose between two alternatives, A and B. The experimenter records both the choice and the time taken to make it. A common approach to modeling this behavior is through a stochastic process $X(t)$ that starts at zero and evolves randomly over time until it first reaches +a{+}a or a{-}a, where $a$ is the evidence threshold. The choice (A vs. B) is modeled by which boundary is hit (+a{+}a vs. a{-}a) and the response time is modeled by the first hitting time.
@@
This basic DDM can be modified in a variety of ways to better model experimental data and test theories about the biological processes underlying decision making. For instance, the boundaries ($\pm a$) can be asymmetric and vary over time to reflect changing pressures to make a decision; the starting position $X(0)$ can be nonzero or random, representing an initial decision bias; the rates of drifting towards the boundaries can depend on the choices and vary over time; and the random process $X(t)$ can have more complicated stochastic dynamics than Brownian motion [14]. We adopt the terminology from [50] and call these models generalized drift diffusion models (GDDMs). Simple, computationally tractable versions of GDDMs have found widespread use in cognitive neuroscience and mathematical psychology [44, 45, 17, 11, 64]. While more complex versions of GDDMs have been proposed to encode a broader range of scientific hypotheses, their quantitative assessment and applications are currently limited due to computational challenges in parameter estimation [14]. In particular, there do not exist fast algorithms for calculating the likelihood function of the parameters (i.e., the probability density of the hitting time on the boundary reached) in some of the more complex GDDMs that are currently of scientific interest.
Computational speed is crucial in many cases for the following reason: As noted in [32], first passage time models commonly incorporate covariates that may vary across individuals and even within a single trial. This is particularly relevant in computational cognitive neuroscience, where the precise specification of the GDDM might change every time a subject makes a choice, i.e., in each experimental trial. An experiment with, say, $50$ subjects, each completing $200$ trials, might require evaluating the likelihoods of $10^{4}$ different versions of the GDDM to get the overall dataset likelihood. For instance, in equation (1) the drift rate $\mu$ might be different on every trial if the choices presented to subjects are different on every trial. The problem becomes even more difficult when the trial-specific drift rate can vary dynamically within a trial.
@@
For a specific example, our original motivation to study this problem stems from attentional drift diffusion models (aDDMs)[29], which have been proposed as a generative model for investigating how visual attention influences decision making. In aDDMs, the drift rate $\mu(t)$ in each trial is allowed to covary on a moment-by-moment basis, depending on the subject’s eye movements toward one choice option or another, which are also recorded in the experimental setup. The entire shape of $\mu(t)$ changes dramatically on every trial. This model is a generic example of the general class of GDDMs where instantaneous within-trial dynamics can be based on exogenous covariates such as brain activity or eye gaze position, hence complicating the numerical procedures for computing the first passage time densities and further applications. On top of this, statistical inference tasks would still require repeated evaluations of the total dataset likelihood at different parameter values, both in frequentist and Bayesian frameworks. In particular, hierarchical Bayesian inference is widely employed in cognitive science to estimate both group- and subject-level parameters simultaneously [63, 14]. The heavy reliance on iterative Markov chain Monte Carlo (MCMC) methods in this setting further intensifies the need for fast likelihood evaluations.
@@
for possibly random $X(0)$ and for functions $\mu$ and $\sigma$ satisfying the so-called Cherkasov condition [7, 47]. Loosely speaking, the Cherkasov condition means that the complicated SDE in equation (2) can be transformed into a simpler SDE more akin to equation (1). A special case that includes many cases of practical interest, such as the aDDM, is when $\mu$ and $\sigma$ depend only on $t$ and not on X(t)X(t) [^1] In the cognitive science context, the terminology “DDM” is typically adopted to denote these models that directly integrate noisy accumulating evidence. The more general model in equation (2), including some examples shown later in this paper (e.g., the Ornstein-Uhlenbeck model), fall within the GDDM framework.. There exist numerical approaches and software for approximating hitting time densities in models even more general than equation (2), but in many cases they are currently too slow for practical use in real-world psychological experiments.
Our approach shares some conceptual similarities to a line of previous work [60, 39, 43, 61, 26]. Both those methods and ours transform the problem into a canonical representation, subdivide time into discrete stages that allow for analytical approximations, compute the final approximation through integration over these stages, and then map the result back to the original problem. However, none of these earlier approaches are directly applicable to our setting or capable of statistical computation at the scale of the present work (see Section 3). In contrast, by leveraging accurate analytical results and numerical integration in a sequential manner, our semi-analytical-numerical approach achieves substantially higher computational efficiency. As shown in Section 5.4.1, it only requires microseconds for our method to evaluate a single likelihood value in a realistic aDDM example. These innovations and efficiency gains are crucial for practical likelihood-based inference in modern experimental settings.
@@
The remainder of this article is structured as follows. Section 2 begins with a mathematical description of the GDDM. Section 3 summarizes previous work related to this paper, including other efforts to address the computational challenges in first passage time models. In Section 4, we introduce our proposed method for computing the FPTDs of certain GDDMs, starting with analytical results for simple DDMs, then combining them sequentially based on Markovian properties. We also apply variable transformations and piecewise linear approximations to extend the method to more complex GDDMs, including those that allow for position-dependent drift and diffusion terms as well as dynamic decision boundaries [57, 14, 19]. Section 5 shows several numerical examples that demonstrate the speed and accuracy of our method for both likelihood computation and statistical inference tasks. Section 6 summarizes the present work, including its applicability and suggestions for future work. It is worth mentioning that while our examples are drawn from cognitive neuroscience and mathematical psychology, the method we propose is equally applicable to first passage time models in other fields. For instance, it could be extended to financial applications such as pricing double barrier options [33], as well as other areas where first passage times play a critical role.
@@ 2 Drift Diffusion Model
where $\mathbbm{1}_{E}(\omega)$ is the indicator function of $E$ that is $1$ if $\omega\in E$ and 0 otherwise.
If we consider the possible truncation of the process at $t=T<\infty$, the process can either exit through the upper boundary $u(t)$ before $T$, exit through the lower boundary $\ell(t)$ before $T$, or fail to hit either boundary by $T$ and instead reach the truncation boundary $t=T$, resulting in three sub-probability densities:
@@
As mentioned in the Introduction, we are motivated by practical problems for which we need to evaluate many different joint densities (3). We define $\boldsymbol{\Xi}=\{\mu(\cdot),\sigma(\cdot),u(\cdot),\ell(\cdot),q_{0}(\cdot)\}$ to be the configuration of a GDDM, where $q_{0}$ is a representation for the distribution of $X(0)$, usually a probability density function or a point mass at a deterministic starting point $x_{0}$. In statistical applications, $\boldsymbol{\Xi}$ will depend on unknown parameters and known covariates, some of which might vary within and across each trial. If the $i$-th trial results in the observation $(\tau_{i},C_{i})$, then the contribution to the likelihood for that trial is $f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i})$. This is simply the FPTD on the boundary $C_{i}$ evaluated at time $\tau_{i}$ using the GDDM configuration $\boldsymbol{\Xi}_{i}$ for the $i$-th trial. In experiments with a time limit, some trials might result in non-response, in which case the contribution to the likelihood is the NPP for that GDDM configuration, say $Q_{i}=Q(\boldsymbol{\Xi}_{i})$. Consequently, denoting $R_{i}=\mathbbm{1}\{\text{a response on trial $i$}\}$, the likelihood function is then given by [^3] The statistical model for $\tau$ here is a mixed distribution consisting of a continuous part supported on $\tau\in(0,T]$ (response) and a point mass at $\tau>T$ (no response). Mixed likelihoods of this form are common in survival analysis, for example see the Section 2.2 of [31].
$\mathcal{L}_{\text{GDDM}}(\boldsymbol{\theta}\mid((\tau_{i},C_{i}))_{i=1}^{n})=$
$\quad\prod_{i=1}^{n}f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i})^{R_{i}}Q(\boldsymbol{\Xi}_{i})^{1-R_{i}}.$
A usual experiment in the cognitive sciences, where GDDMs are ubiquitously applied, produces a dataset with thousands or tens of thousands of trials (across multiple subjects). Consequently, fast and accurate methods for evaluating FPTDs and NPPs are crucial for the practical use of likelihood-based inference in these models.
@@ 3 Related Work
Analytic or semi-analytic methods. For certain simple DDMs, analytic formulas for the first passage time densities or distributions can be derived, often in the form of infinite series. These methods are fast and accurate, but are typically applicable only for very specific cases. For example, they are available for the standard DDM where all parameters $\mu,\sigma,x_{0},u,\ell$ are constant [38, 4, 21]. More recently, [53] used martingale theory to derive closed-form formulas that extend certain theoretical analyses to GDDMs.
The approach in this paper is based on numerically extending these fast special cases to much broader classes of GDDMs. Along this line, earlier works [60, 39, 43, 61] computed NPPs, while [26], [25], and the present work target FPTDs, although [26] and [25] restrict attention to the special case of a one-sided boundary. The FPTD case is more challenging but also more useful for statistical inference since it directly gives the likelihood. A key limitation of many prior approaches [60, 43, 61, 25, 26], is that they rely on Monte Carlo integration, which is infeasible for large-scale real-world applications.
@@
Numerical solutions of governing equations. These methods exploit the known connections between diffusion processes and partial differential equations (PDEs), called Kolmogorov equations (see the Chapter 8 of [62] for an introduction, and Chapter 5.7 of [27] for a mathematically rigorous treatment). Both the Kolmogorov forward equation (KFE, also known as Fokker-Planck equation in physics) [50, 10] and the Kolmogorov backward equation (KBE) [59, 58] have been exploited to derive the FPTDs. Moreover, the first passage time problem can be formulated into Volterra integral equations of renewal type [51, 42], where applications to DDMs [52, 40] have also been explored. In these works, numerical methods, particularly finite difference methods [55], have been developed to solve these equations, enabling the efficient computation of FPTDs in certain cases. Other numerical methods like discontinuous Galerkin methods have also been considered, see [41] and references therein. However, these approaches remain computationally expensive, especially in settings where the GDDM configuration differs across trials, including but not limited to cases with trial-specific covariates that vary dynamically over time. For a comprehensive review of these methods, we refer the readers to [48].
Stochastic simulation. Methods that can sample from the hitting time distribution of the target GDDM can be used to approximate FPTDs via density estimation. Efficient methods for exact sampling are challenging; see [23, 24] and the references therein. If the goal is statistical inference, then these sampling algorithms can be incorporated into a variety of simulation-based inference (SBI) procedures; see [8] for review and [13] for an SBI example with DDMs. Nevertheless, the Monte Carlo methods underlying SBI are computationally intensive and can become impractical for large datasets, particularly when $\Xi$ involves high-dimensional parameters or covariates that vary dynamically over time and across trials.
@@ 4.1 Single-stage Model
[22] derives the following analytical formulas [^4] Another work along this direction is [1]. We adopt the results in [22] since they are simpler and were shown to agree with those of [1]. Equation (4.1), (4.1), and (4.1) have been slightly rewritten for use in this paper, compared to their original form in [22]. for FPTDs and NPDs in the single-stage model for the special case when $x_{0}=0$ and $\sigma=1$: We define $\overline{a}=\frac{a_{1}+a_{2}}{2},\ \overline{b}=\frac{b_{1}+b_{2}}{2},\ c=a_{1}-a_{2},\ b=\frac{b_{2}-b_{1}}{2}$, then
@@
    $\quad\frac{1}{\sqrt{2\pi t^{3}}}e^{-\frac{b}{c}a_{1}^{2}+a_{1}\delta_{u}-\frac{1}{2}\delta_{u}^{2}t}$
    $\quad\sum_{j=0}^{\infty}(-1)^{j}\alpha_{j}e^{\left(\frac{b}{c}-\frac{1}{2t}\right)\alpha_{j}^{2}}\mathbbm{1}_{(0,T]}(t)$
@@
    $\quad\frac{1}{\sqrt{2\pi t^{3}}}e^{-\frac{b}{c}a_{2}^{2}-a_{2}\delta_{\ell}-\frac{1}{2}\delta_{\ell}^{2}t}$
    $\quad\sum_{j=0}^{\infty}(-1)^{j}\beta_{j}e^{\left(\frac{b}{c}-\frac{1}{2t}\right)\beta_{j}^{2}}\mathbbm{1}_{(0,T]}(t)$
@@
    $\quad\frac{1}{\sqrt{2\pi T}}e^{(\mu-\overline{b})x-\frac{1}{2}(\mu^{2}-\overline{b}^{2})T}\bigg\{e^{-\frac{y^{2}}{2T}}+$
    $\quad\sum_{j=1}^{\infty}\bigg[e^{4bj(jc-\overline{a})-\frac{(y-2jc)^{2}}{2T}}$-
    $\quad e^{2b(2j-1)\left(jc-a_{1}\right)-\frac{(y+2jc-2a_{1})^{2}}{2T}}+$
    $\quad e^{4bj(jc+\overline{a})-\frac{(y+2jc)^{2}}{2T}}$-
    $\quad e^{2b(2j-1)\left(jc+a_{2}\right)-\frac{(y-2jc-2a_{2})^{2}}{2T}}\bigg]\bigg\}$
    $\quad\mathbbm{1}_{(\ell(T),u(T))}(x).$
@@
$f_{u}^{\text{single}}(t;\mu,\sigma,\mathcal{B},x_{0})=$
$\quad f_{u}^{\text{basic}}\Big(t;\tfrac{\mu}{\sigma},\tfrac{a_{1}-x_{0}}{\sigma},\tfrac{b_{1}}{\sigma},\tfrac{a_{2}-x_{0}}{\sigma},\tfrac{b_{2}}{\sigma},T\Big),$ (9)
$f_{\ell}^{\text{single}}(t;\mu,\sigma,\mathcal{B},x_{0})=$
$\quad f_{\ell}^{\text{basic}}\Big(t;\tfrac{\mu}{\sigma},\tfrac{a_{1}-x_{0}}{\sigma},\tfrac{b_{1}}{\sigma},\tfrac{a_{2}-x_{0}}{\sigma},\tfrac{b_{2}}{\sigma},T\Big),$ (10)
$q^{\text{single}}(x;\mu,\sigma,\mathcal{B},x_{0})=$
$\quad\tfrac{1}{\sigma}q^{\text{basic}}\Big(\tfrac{x-x_{0}}{\sigma};\tfrac{\mu}{\sigma},\tfrac{a_{1}-x_{0}}{\sigma},\tfrac{b_{1}}{\sigma},\tfrac{a_{2}-x_{0}}{\sigma},\tfrac{b_{2}}{\sigma},T\Big).$ (11)
@@
Figure 2: (a) A single-stage model is a GDDM with constant drift $\mu$ and two linear boundaries $u(t),\ell(t)$ starting at $x=x_{0}$ and truncated at $t=T$. The process can hit one of $u(t)$ and $\ell(t)$ before $T$, or hit neither. This results in three distinct sub-probability distributions, corresponding to these three outcomes, with their sub-probability densities plotted by red, purple, and green curves along the boundaries. Sample trajectories, colored to match their respective outcomes, are also displayed. The blue line represents the expected value of the process, given by $\mathbb{E}[X(t)]=x_{0}+\mu t$. In a single-stage model the total area under the three sub-probability density curves sums to 1. (b) The integrated model extends the single-stage model by introducing random variability in the initial position $X(0)$. The expected value of the process in this model is given by $\mathbb{E}[X(t)]=\mathbb{E}[X(0)]+\mu t$, represented by the blue line. The density of $X(0)$, denoted by $q_{0}$, is illustrated by the grey curve. We allow $q_{0}$ to be a sub-probability density function. In an integrated single-stage model, the total area under the three outcome sub-probability density curves sums to the area under the initial grey curve.